OpenStudy (anonymous):

prove that value of sqrt 100 - sqrt 99 lies vetween 1/18 and 1/20

OpenStudy (king):

we can say that sqrt99=3sqrt11=3*3.316=9.948 sqrt100=10 so 10-9.948=0.052 nw 1/18=0.55555... 1/20=0.5 so 0.052 lies between 0.55555... and 0.5 Thus,Proved!!!!!!

OpenStudy (experimentx):

10 - sqrt(100-1) => 10 - 10(1 - 1/100)^1/2 now let's expand (1 - 1/100)^1/2 using binomial theorem

OpenStudy (king):

Hello guest please sign in or sign up...Thnk u!!:D

OpenStudy (king):

@experimentX u can use binomial theorem but then hw do we know if coolbird143 knows this or not?

OpenStudy (experimentx):

(1-1/100)^1/2 = 1 -1/100x1/2-(1/100)^2x1/2 - ...

OpenStudy (experimentx):

well the best way is to use calculator ... :D

OpenStudy (king):

or use ure brains...using calculator is bad....

OpenStudy (experimentx):

10 - 10(1 - 1/100)^1/2 = 10 - 10 + 1/2*10/100 + 1/2*1/2*1/2*10/10000 + 1/2*1/2*3/2*1/6*10/1000000 => 1/2*10/100 + 1/2*1/2*1/2*10/10000 + 1/2*1/2*3/2*1/6*10/1000000 ... which is a converging series that lie between 1/18 and 1/20